\begin{thebibliography}{35}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Addawood and Bashir(2016)}]{addawood:16}
Aseel Addawood and Masooda Bashir. 2016.
\newblock "what is your evidence?" a study of controversial topics on social
  media.
\newblock In \emph{Proceedings of the Third Workshop on Argument Mining
  (ArgMining2016)}, pages 1--11.

\bibitem[{Applebee et~al.(2003)Applebee, Langer, Nystrand, and
  Gamoran}]{Applebee:03}
Arthur~N Applebee, Judith~A Langer, Martin Nystrand, and Adam Gamoran. 2003.
\newblock Discussion-based approaches to developing understanding: Classroom
  instruction and student performance in middle and high school english.
\newblock \emph{American Educational Research Journal}, 40(3):685--730.

\bibitem[{Beach and Myers(2001)}]{Beach:01}
Richard Beach and Jamie Myers. 2001.
\newblock \emph{Inquiry-based English instruction: Engaging students in life
  and literature}, volume~55.
\newblock Teachers College Press.

\bibitem[{Bloome et~al.(2005)Bloome, Carter, Christian, Otto, and
  Shuart-Faris}]{Bloome:05}
David Bloome, Stephanie~Power Carter, Beth~Morton Christian, Sheila Otto, and
  Nora Shuart-Faris. 2005.
\newblock \emph{Discourse analysis and the study of classroom language and
  literacy events: A microethnographic perspective}.
\newblock Lawrence Erlbaum.

\bibitem[{Chisholm and Godley(2011)}]{Chisholm:11}
James~S Chisholm and Amanda~J Godley. 2011.
\newblock Learning about language through inquiry-based discussion: Three
  bidialectal high school students’ talk about dialect variation, identity,
  and power.
\newblock \emph{Journal of Literacy Research}, 43(4):430--468.

\bibitem[{Danielson(2011)}]{Danielson:11}
Charlotte Danielson. 2011.
\newblock Evaluations that help teachers learn.
\newblock \emph{Educational leadership}, 68(4):35--39.

\bibitem[{Elizabeth et~al.(2012)Elizabeth, Ross~Anderson, Snow, and
  Selman}]{Elizabeth:12}
Tracy Elizabeth, Trisha~L Ross~Anderson, Elana~H Snow, and Robert~L Selman.
  2012.
\newblock Academic discussions: An analysis of instructional discourse and an
  argument for an integrative assessment framework.
\newblock \emph{American Educational Research Journal}, 49(6):1214--1250.

\bibitem[{Engle and Conant(2002)}]{Engle:02}
Randi~A Engle and Faith~R Conant. 2002.
\newblock Guiding principles for fostering productive disciplinary engagement:
  Explaining an emergent argument in a community of learners classroom.
\newblock \emph{Cognition and Instruction}, 20(4):399--483.

\bibitem[{Ghosh et~al.(2016)Ghosh, Khanam, Han, and Muresan}]{Ghosh:16}
Debanjan Ghosh, Aquila Khanam, Yubo Han, and Smaranda Muresan. 2016.
\newblock Coarse-grained argumentation features for scoring persuasive essays.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, volume~2, pages
  549--554.

\bibitem[{Godley and Olshefski(2017)}]{Godley:17}
Amanda Godley and Christopher Olshefski. 2017.
\newblock The role of argument moves, specificity and evidence type in
  meaningful literary discussions across diverse secondary classrooms.
\newblock Unpublished paper presented at Literacy Research Association 67th
  Annual Conference: Literacy Research for Expanding Meaningfulness.

\bibitem[{Grossman et~al.(2014)Grossman, Cohen, Ronfeldt, and
  Brown}]{Grossman:14}
Pam Grossman, Julie Cohen, Matthew Ronfeldt, and Lindsay Brown. 2014.
\newblock The test matters: The relationship between classroom observation
  scores and teacher value added on multiple types of assessment.
\newblock \emph{Educational Researcher}, 43(6):293--303.

\bibitem[{Kim(2014)}]{Kim:14}
Il-Hee Kim. 2014.
\newblock Development of reasoning skills through participation in
  collaborative synchronous online discussions.
\newblock \emph{Interactive Learning Environments}, 22(4):467--484.

\bibitem[{Klebanov et~al.(2016)Klebanov, Stab, Burstein, Song, Gyawali, and
  Gurevych}]{Klebanov:16}
Beata~Beigman Klebanov, Christian Stab, Jill Burstein, Yi~Song, Binod Gyawali,
  and Iryna Gurevych. 2016.
\newblock Argumentation: Content, structure, and relationship with essay
  quality.
\newblock In \emph{Proceedings of the Third Workshop on Argument Mining
  (ArgMining2016)}, pages 70--75.

\bibitem[{Lee(2006)}]{Lee:06}
Carol~D Lee. 2006.
\newblock ‘every good-bye ain’t gone’: analyzing the cultural
  underpinnings of classroom talk.
\newblock \emph{International Journal of Qualitative Studies in Education},
  19(3):305--327.

\bibitem[{Li and Nenkova(2015)}]{Li:15}
Junyi~Jessy Li and Ani Nenkova. 2015.
\newblock Fast and accurate prediction of sentence specificity.
\newblock In \emph{Proceedings of the Twenty-Ninth Conference on Artificial
  Intelligence (AAAI)}, pages 2281--2287.

\bibitem[{Li et~al.(2016)Li, O’Daniel, Wu, Zhao, and Nenkova}]{Li:16}
Junyi~Jessy Li, Bridget O’Daniel, Yi~Wu, Wenli Zhao, and Ani Nenkova. 2016.
\newblock Improving the annotation of sentence specificity.
\newblock In \emph{Proceedings of the Tenth International Conference on
  Language Resources and Evaluation (LREC)}.

\bibitem[{Louis and Nenkova(2011)}]{Louis:11}
Annie Louis and Ani Nenkova. 2011.
\newblock General versus specific sentences: automatic identification and
  application to analysis of news summaries.
\newblock Technical Report MS-CIS-11-07, University of Pennsylvania.

\bibitem[{Louis and Nenkova(2012)}]{Louis:12}
Annie Louis and Ani Nenkova. 2012.
\newblock A corpus of general and specific sentences from news.
\newblock In \emph{LREC}, pages 1818--1821.

\bibitem[{Lugini and Litman(2017)}]{Lugini:17}
Luca Lugini and Diane Litman. 2017.
\newblock Predicting specificity in classroom discussion.
\newblock In \emph{Proceedings of the 12th Workshop on Innovative Use of NLP
  for Building Educational Applications}, pages 52--61.

\bibitem[{McHugh(2012)}]{Mchugh:12}
Mary~L McHugh. 2012.
\newblock Interrater reliability: the kappa statistic.
\newblock \emph{Biochemia medica: Biochemia medica}, 22(3):276--282.

\bibitem[{Michaels et~al.(2008)Michaels, O’Connor, and Resnick}]{Michaels:08}
Sarah Michaels, Catherine O’Connor, and Lauren~B Resnick. 2008.
\newblock Deliberative discourse idealized and realized: Accountable talk in
  the classroom and in civic life.
\newblock \emph{Studies in philosophy and education}, 27(4):283--297.

\bibitem[{Murphy et~al.(2009)Murphy, Wilkinson, Soter, Hennessey, and
  Alexander}]{Murphy:09}
P~Karen Murphy, Ian~AG Wilkinson, Anna~O Soter, Maeghan~N Hennessey, and John~F
  Alexander. 2009.
\newblock Examining the effects of classroom discussion on students’
  comprehension of text: A meta-analysis.
\newblock \emph{Journal of Educational Psychology}, 101(3):740.

\bibitem[{Nguyen and Litman(2018)}]{Nguyen:18}
Huy~V Nguyen and Diane~J Litman. 2018.
\newblock Argument mining for improving the automated scoring of persuasive
  essays.
\newblock In \emph{Proceedings of the Thirty-Second AAAI Conference on
  Artificial Intelligence}.

\bibitem[{Niculae et~al.(2017)Niculae, Park, and Cardie}]{Niculae:17}
Vlad Niculae, Joonsuk Park, and Claire Cardie. 2017.
\newblock {Argument Mining with Structured SVMs and RNNs}.
\newblock In \emph{Proceedings of ACL}.

\bibitem[{Peldszus and Stede(2013)}]{Peldszus:13}
Andreas Peldszus and Manfred Stede. 2013.
\newblock From argument diagrams to argumentation mining in texts: A survey.
\newblock \emph{International Journal of Cognitive Informatics and Natural
  Intelligence (IJCINI)}, 7(1):1--31.

\bibitem[{Peldszus and Stede(2015)}]{Peldszus:15}
Andreas Peldszus and Manfred Stede. 2015.
\newblock Joint prediction in mst-style discourse parsing for argumentation
  mining.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing}, pages 938--948.

\bibitem[{Persing and Ng(2016)}]{Persing:16}
Isaac Persing and Vincent Ng. 2016.
\newblock End-to-end argumentation mining in student essays.
\newblock In \emph{Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 1384--1394.

\bibitem[{Resnick and Schantz(2015)}]{Resnick:15}
Lauren~B Resnick and Faith Schantz. 2015.
\newblock Talking to learn: The promise and challenge of dialogic teaching.
\newblock \emph{Socializing Intelligence through academic talk and dialogue},
  pages 441--450.

\bibitem[{Reznitskaya and Gregory(2013)}]{Reznitskaya:13}
Alina Reznitskaya and Maughn Gregory. 2013.
\newblock Student thought and classroom language: Examining the mechanisms of
  change in dialogic teaching.
\newblock \emph{Educational Psychologist}, 48(2):114--133.

\bibitem[{Reznitskaya et~al.(2009)Reznitskaya, Kuo, Clark, Miller, Jadallah,
  Anderson, and Nguyen-Jahiel}]{Reznitskaya:09}
Alina Reznitskaya, Li-Jen Kuo, Ann-Marie Clark, Brian Miller, May Jadallah,
  Richard~C Anderson, and Kim Nguyen-Jahiel. 2009.
\newblock Collaborative reasoning: A dialogic approach to group discussions.
\newblock \emph{Cambridge Journal of Education}, 39(1):29--48.

\bibitem[{Sohmer et~al.(2009)Sohmer, Michaels, O’Connor, and
  Resnick}]{Sohmer:09}
Richard Sohmer, Sarah Michaels, MC~O’Connor, and Lauren Resnick. 2009.
\newblock Guided construction of knowledge in the classroom.
\newblock \emph{Transformation of knowledge through classroom interaction},
  pages 105--129.

\bibitem[{Soter et~al.(2008)Soter, Wilkinson, Murphy, Rudge, Reninger, and
  Edwards}]{Soter:08}
Anna~O Soter, Ian~A Wilkinson, P~Karen Murphy, Lucila Rudge, Kristin Reninger,
  and Margaret Edwards. 2008.
\newblock What the discourse tells us: Talk and indicators of high-level
  comprehension.
\newblock \emph{International Journal of Educational Research}, 47(6):372--391.

\bibitem[{Stab and Gurevych(2017)}]{Stab:17}
Christian Stab and Iryna Gurevych. 2017.
\newblock Parsing argumentation structures in persuasive essays.
\newblock \emph{Computational Linguistics}, 43(3):619--659.

\bibitem[{Toulmin(1958)}]{Toulmin:58}
Stephen Toulmin. 1958.
\newblock \emph{The uses of argument}.
\newblock Cambridge: Cambridge University Press.

\bibitem[{Wachsmuth et~al.(2016)Wachsmuth, Al~Khatib, and Stein}]{Wachsmuth:16}
Henning Wachsmuth, Khalid Al~Khatib, and Benno Stein. 2016.
\newblock Using argument mining to assess the argumentation quality of essays.
\newblock In \emph{Proceedings of COLING 2016, the 26th International
  Conference on Computational Linguistics: Technical Papers}, pages 1680--1691.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{naaclhlt2018}
\usepackage{times}
\usepackage{latexsym}
\usepackage{multirow}
\usepackage{graphicx}

\usepackage{url}

\aclfinalcopy 

\title{Annotating Student Talk in Text-based Classroom Discussions}

\author{Luca Lugini, Diane Litman, Amanda Godley, Christopher Olshefski\\
  University of Pittsburgh \\
  Pittsburgh, PA 15260 \\
  {\tt \{lul32,dlitman,agodley,cao48\}@pitt.edu}}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Classroom discussions in English Language Arts have a positive effect on students' reading, writing, and reasoning skills.
Although prior work has largely focused on teacher talk and student-teacher interactions, 
we focus on three theoretically-motivated aspects of high-quality student talk: argumentation, specificity, and knowledge domain.
We introduce an annotation scheme, then show that the scheme can be used to produce reliable annotations and that the annotations are predictive of discussion quality.
We also highlight opportunities provided by our scheme for educational and natural language processing research.
\end{abstract}

\section{Introduction}
Current research, theory, and policy surrounding K-12 instruction in the United States highlight the role of student-centered disciplinary discussions
(i.e. discussions related to a specific academic discipline or school subject such as physics or English Language Arts)
in instructional quality and student learning opportunities \cite{Danielson:11,Grossman:14}. Such student-centered discussions -- often called ``dialogic" or ``inquiry-based'' -- are widely viewed as the most effective instructional approach for disciplinary understanding, problem-solving, and literacy \cite{Elizabeth:12,Engle:02, Murphy:09}. In English Language Arts (ELA) classrooms, student-centered discussions about literature have a positive impact on the development of students' reasoning, writing, and reading skills \cite{Applebee:03, Reznitskaya:13}. However, most studies have focused on the role of teachers and their talk \cite{Bloome:05,Elizabeth:12,Michaels:08} rather than on the aspects of student talk that contribute to discussion quality. 

Additionally, studies of student-centered discussions rarely use the same coding schemes, making it difficult to generalize across studies \cite{Elizabeth:12,Soter:08}. This limitation is partly due to the time-intensive work required to analyze discourse data through qualitative methods such as ethnography and discourse analysis. Thus, qualitative case studies have generated compelling theories about the specific features of student talk that lead to high-quality discussions, but few findings
can be generalized and leveraged to influence instructional improvements across ELA classrooms. 

As a first step towards developing an automated system for detecting the features of student talk that lead to high quality discussions, we propose a new annotation scheme for student talk during ELA ``text-based" discussions - that is, discussions that center on a text or piece of literature (e.g., book, play, or speech).
The annotation scheme was developed to capture three aspects of classroom talk that are theorized in the literature as important to discussion quality and learning opportunities: {\it argumentation} (the process of systematically reasoning in support of an idea),
{\it specificity} (the quality of belonging or relating uniquely to a particular subject),
and {\it knowledge domain} (area of expertise represented in the content of the talk).
We demonstrate the reliability and validity of our scheme via an annotation study of five transcripts of classroom discussion.


\section{Related Work}
One discourse feature used to assess the quality of discussions is students' argument moves: their claims about the text, their sharing of textual evidence for claims, and their warranting or reasoning to support the claims \cite{Reznitskaya:09,Toulmin:58}. Many researchers view student reasoning as of primary importance, particularly when the reasoning is elaborated and highly inferential \cite{Kim:14}. In Natural Language Processing (NLP), most educationally-oriented argumentation research 
has focused on corpora of student persuasive essays \cite{Ghosh:16,Klebanov:16,Persing:16,Wachsmuth:16,Stab:17,Nguyen:18}.
We instead focus on 
multi-party spoken discussion transcripts from classrooms. A second key difference consists in the inclusion of the warrant label in our scheme, as it is important to understand how students explicitly use reasoning to connect evidence to claims. 

Educational studies suggest that discussion quality is also influenced by the specificity of student talk \cite{Chisholm:11,Sohmer:09}. Chisholm and Godley found that as specificity increased, the quality of students' claims and reasoning also increased. Previous NLP research has studied specificity in the context of professionally written newspaper articles
\cite{Li:15,Li:16,Louis:11,Louis:12}.
While the annotation instructions used in these studies work well for general purpose corpora,
specificity in text-based discussions 
also needs to capture particular relations between discussions and texts. Furthermore, since the concept of a sentence is not clearly defined in speech,
we annotate argumentative discourse units rather than sentences  (see Section~\ref{sec:scheme}). 

The knowledge domain of student talk may also matter, that is, whether the talk focuses on disciplinary knowledge or lived experiences. Some research suggests that disciplinary learning opportunities are maximized when students draw on evidence and reasoning that are commonly accepted in the discipline \cite{Resnick:15}, although some studies suggest that evidence or reasoning from lived experiences increases discussion quality \cite{Beach:01}.
Previous related work in NLP analyzed evidence type for argumentative tweets \cite{addawood:16}. Although the categories of evidence type are different, their definition of evidence type is in line with our definition of knowledge domain. However, our research is distinct from this research in its application domain (i.e. social media vs. education) and in analyzing knowledge domain for all argumentative components, not only those containing claims.

\section{Annotation Scheme}
\label{sec:scheme}
Our annotation scheme\footnote{The  coding manual is in the supplemental material.} uses argument moves as the unit of analysis. We define an argument move as an utterance, or part of an utterance, that contains an argumentative discourse unit (ADU) \cite{Peldszus:13}.
Like Peldszus and Stede \shortcite{Peldszus:15}, in this paper we use  transcripts already segmented into argument moves and focus on the steps following segmentation, i.e., labeling argumentation, specificity, and knowledge domain.
Table \ref{tab:examples} shows a section of a transcribed classroom discussion along with labels assigned by a human annotator following segmentation.
\begin{table*}[ht]
\centering
\begin{tabular}{|c|c|p{0.425\linewidth}|c|c|c|}
\hline
\textbf{Move} & \textbf{Stu} & \textbf{Argument Move} & \textbf{Argument} & \textbf{Specificity} & \textbf{Domain}\\ \hline
23 & S1 & She's like really just protecting Willy from everything.  & claim & medium   & disciplinary  \\ \hline
24 & S1 & Like at the end of the book remember how she was telling the kids to leave and never come back. & evidence  & medium & disciplinary     \\ \hline
25 & S1 & Like she's not even caring about them, she's caring about Willy.  & warrant & medium & disciplinary   \\ \hline
\multicolumn{1}{c}{} \\[-0.15cm]
\hline
41 & S2 & It's like she's concerned with him trying to [inaudible] and he's concerned with trying to make her happy, you know? So he feels like he's failing when he's not making her happy like  & claim & high  & disciplinary \\ \hline
42 & S2 & "Let's bring your mother some good news"  & evidence  & high  & disciplinary\\ \hline
43 & S2 & but she knew that, there wasn't any good news,  so she wanted to act happy so he wouldn't be in pain. & warrant & high  & disciplinary \\ \hline
\multicolumn{1}{c}{} \\[-0.15cm]
\hline
55 & S3 & Some people they just ask for a job is just like, some money. & evidence  & low & experiential \\ \hline
\end{tabular}
\caption{Examples of argument moves and their respective annotations from a discussion of the book \textit{Death of a Salesman}. As shown by the argument move numbers, boxes for students S1, S2, and S3 indicate separate, non contiguous excerpts of the discussion.}
\label{tab:examples}
\end{table*}

\subsection{Argumentation}
The argumentation scheme is based on \cite{Lee:06} and consists of a simplified set of labels derived from Toulmin's \shortcite{Toulmin:58} model:
$(i)$ {\it Claim}: an arguable statement that presents a particular interpretation of a text or topic.
$(ii)$ {\it Evidence}: facts, documentation, text reference, or testimony used to support or justify a claim.
$(iii)$ {\it Warrant}: reasons explaining how a specific evidence instance supports a specific claim.
Our scheme specifies that warrants must come after claim and evidence, since by definition warrants cannot exist without them.

The first three moves in Table \ref{tab:examples} show a natural expression of an argument: a student first claims that Willy's wife is only trying to protect him, then provides a reference as evidence by mentioning something she said to her kids at the end of the book, and finally explains how
not caring
about her kids ties the evidence to the initial claim.
The second group shows the same argument progression, with evidence given as a direct quote.

\subsection{Specificity}
Specificity annotations are based on \cite{Chisholm:11} and have the goal of capturing text-related characteristics expressed in student talk.
Specificity labels are directly related to four distinct elements for an argument move:
(1) it is specific to one (or a few) character or scene;
(2) it makes significant qualifications or elaborations;
(3) it uses content-specific vocabulary (e.g. quotes from the text);
(4) it provides a chain of reasons.
Our annotation scheme for specificity includes three labels along a linear scale:
$(i)$ {\it Low}: statement that does not contain any of these elements.
$(ii)$ {\it Medium}: statement that accomplishes one of these elements.
$(iii)$ {\it High}: statement that clearly accomplishes at least two specificity elements.
Even though we do not explicitly use labels for the four specificity elements, we found that explicitly breaking down specificity into multiple components helped increase reliability when training annotators.

The first three argument moves in Table \ref{tab:examples} all contain the first element, as they refer to select characters in the book. However, no content-specific vocabulary,  clear chain of reasoning, or significant qualifications are provided; therefore all three moves are labeled as medium specificity.
The fourth move, however, accomplishes the first and fourth specificity elements, and is labeled as  high specificity.
The fifth move is also labeled high specificity since it is specific to one character/scene, and provides a direct quote from the text.
The last  move is labeled as low specificity as it reflects an overgeneralization about all humans.

\subsection{Knowledge Domain}

The possible labels for knowledge domain are:
$(i)$ {\it Disciplinary}: the statement is grounded in knowledge gathered from a text (either the one under discussion or others), such as a quote or a description of a character/event. 
$(ii)$ {\it Experiential}: the statement is drawn from human experience, such as what the speaker has experienced or thinks that other humans have experienced.

In Table \ref{tab:examples} the first six argument moves are labeled as disciplinary, since the moves reflect knowledge from the text currently being discussed. The last move, however, draws from a student's experience or perceived knowledge about the real world.

\section{Reliability and Validity Analyses}
\label{sec:data}
We carried out a reliability study for the proposed scheme using two pairs of expert annotators, P1 and P2. The annotators were trained by coding one transcript at a time and discussing disagreements.
Five text-based discussions were used for testing reliability after training:
pair P1 annotated discussions of \textit{The Bluest Eye}, \textit{Death of a Salesman}, and \textit{Macbeth}, while pair P2 annotated two separate discussions of \textit{Ain't I a Woman}.
250 argument moves (discussed by over 40 students and consisting of over 8200 words) were annotated. Inter-rater reliability was assessed using Cohen's kappa: unweighted  for argumentation and knowledge domain, but quadratic-weighted
 for specificity given its ordered labels.

\begin{table}[t]
\centering
\begin{tabular}{|c|c|c|c|}
\hline
\textbf{Moves} & \textbf{\begin{tabular}[c]{@{}c@{}}Argumen-\\ tation\\ (kappa)\end{tabular}} & \textbf{\begin{tabular}[c]{@{}c@{}}Specificity\\ (qwkappa)\end{tabular}} & \textbf{\begin{tabular}[c]{@{}c@{}}Domain\\ (kappa)\end{tabular}} \\ \hline
169            & 0.729                                                                        & 0.874                                                                    & 0.980                                                             \\ \hline
81             & 0.725                                                                        & 0.930                                                                    & 1                                                                 \\ \hline
\end{tabular}
\caption{Inter-rater reliability for pairs P1 and P2.}
\label{tab:irr}
\end{table}
Table \ref{tab:irr} shows that kappa for argumentation ranges from $0.61 - 0.8$, which generally indicates substantial agreement \cite{Mchugh:12}. Kappa values for specificity and knowledge domain are in the $0.81 - 1$ range which generally indicates almost perfect agreement \cite{Mchugh:12}.
These results show that our proposed annotation scheme can be used to produce reliable annotations of classroom discussion with respect to argumentation, specificity, and knowledge domain.

\begin{table}[t]
\centering
\begin{tabular}{|r|c|c|c|}
\hline
\multicolumn{1}{|l|}{\textbf{Argumentation}}                                              & \textbf{evidence} & \textbf{warrant} & \textbf{claim} \\ \hline
\textbf{evidence}                                                                         & 25                & 5                & 0              \\ \hline
\textbf{warrant}                                                                          & 6                 & 92               & 12             \\ \hline
\textbf{claim}                                                                            & 0                 & 2                & 27             \\ \hline \hline
\multicolumn{1}{|l|}{\textbf{Specificity}}                                                & \textbf{low}      & \textbf{medium}     & \textbf{high}  \\ \hline
\textbf{low}                                                                              & 59                & 5                & 3              \\ \hline
\textbf{medium}                                                                              & 5                 & 25               & 2              \\ \hline
\textbf{high}                                                                             & 1                 & 6                & 63             \\ \hline \hline
\multicolumn{1}{|l|}{\textbf{\begin{tabular}[c]{@{}l@{}}Knowledge\\ Domain\end{tabular}}} & \multicolumn{1}{|c|}{\textbf{\begin{tabular}[c]{@{}l@{}}discipl-\\ inary\end{tabular}}}  & \multicolumn{1}{|c|}{\textbf{\begin{tabular}[c]{@{}l@{}}experi-\\ ential\end{tabular}}}   &                \\ \hline
\textbf{disciplinary}                                                                          & 138               & 1                &                \\ \hline
\textbf{experiential}                                                                            & 0                 & 30               &                \\ \hline
\end{tabular}
\caption{Confusion matrices for argumentation, specificity, and knowledge domain, for annotator pair P1.}
\label{tab:confusion_matrices}
\end{table}

Table \ref{tab:confusion_matrices} shows confusion matrices\footnote{The class distributions for argumentation and specificity labels vary significantly across transcripts, as can be seen in \cite{Lugini:17} and \cite{Godley:17}.}
for annotator pair P1 (we observed similar trends for P2). The argumentation section of the table shows that the largest number of disagreements happens between the claim and warrant labels. One reason may be related to the constraint we impose on warrants - they require the existence of a claim and evidence. If a student tries to provide a warrant for a claim that happened much earlier in the discussion, the annotators might interpret the warrant as new claim.
The specificity section shows relatively few low-high label disagreements as compared to low-med and med-high.
This is also reflected in the quadratic-weighted kappa as low-high disagreements will carry a larger penalty (unweighted kappa is $0.797$).
The main reasons for disagreements over specificity labels come from two of the four specificity elements discussed in Section 3.2: whether an argument move is related to one character or scene, and whether it provides a chain of reasons. With respect to the first of these two elements we observed disagreements in argument moves containing pronouns with an ambiguous reference. Of particular note is the pronoun \textit{it}. If we consider the argument move \textit{``I mean even if you know you have a hatred towards a standard or whatever, you still don't kill it"}, the pronoun \textit{it} clearly refers to something within the move (i.e. the standard) that the student themselves mentioned. In contrast, for argument moves such as \textit{``It did happen"} it might not be clear to what previous move the pronoun refers, therefore creating confusion on whether this specificity element is accomplished.
Regarding specificity element (4) we found that it was easier to determine the presence of a chain of reasons when discourse connectives (e.g. because, therefore) were present in the argument move. The absence of explicit discourse connectives in an argument move might drive annotators to disagree on the presence/absence of a chain of reasons, which is likely to result in a different specificity label.
Additionally, annotators found that shorter turns at talk proved harder to annotate for specificity.
Finally, as we can see from the third section in the table, knowledge domain has the lowest disagreements with only one.

We also \cite{Godley:17} explored the validity of our coding scheme by comparing our annotations of student talk to English Education experts' evaluations (quadratic-weighted kappa of 0.544) of the discussion's quality. Using stepwise regressions, we found that the best model of discussion quality (R-squared of $0.432$) included all three of our coding dimensions: argumentation, specificity, and knowledge domain.

\section{Opportunities and Challenges}
Our annotation scheme introduces opportunities for the educational community to conduct further research on the relationship between features of student talk, student learning, and discussion quality.
Although Chisholm and Godley \shortcite{Chisholm:11} and we found relations between our coding constructs and discussion quality, 
these were small-scale studies based on manual annotations. Once automated classifiers are developed, such relations between talk and learning can be examined at scale. Also, automatic labeling via a standard coding scheme can support the generalization of findings across studies, and potentially lead to
automated tools for teachers and students.

The proposed annotation scheme also introduces NLP opportunities and challenges.
Existing systems for classifying specificity and argumentation have largely been designed to analyze written text rather than spoken discussions.
This is (at least in part) due to a lack of publicly available corpora and schemes for annotating argumentation and specificity in spoken discussions. The development of an annotation scheme explicitly designed for this problem is the first step towards collecting and annotating corpora that can be used by the NLP community to advance the field in this particular area.
Furthermore, in text-based discussions, NLP methods need to tightly couple the discussion with contextual information (i.e., the text under discussion).
For example, an argument move from one of the discussions mentioned in Section 4 stated \textit{``She's saying like free like,  I don't have to be, I don't have to be this salesman's wife anymore, your know? I don't have to play this role anymore."} The use of the term \textit{salesman} shows the presence of specificity element (3) (see Section 3.2) because the text under discussion is indeed \textit{Death of a Salesman}. If the students were discussing another text, the mention of the term \textit{salesman} would not indicate one of the specificity elements, therefore lowering the specificity rating.
Thus, using existing systems is unlikely to yield good performance.  In fact, we previously \cite{Lugini:17} showed that while using an off-the-shelf system for predicting specificity in newspaper articles resulted in low performance when applied to classroom discussions, 
exploiting characteristics of our data could significantly improve performance.
We have similarly evaluated the performance of two existing argument mining systems \cite{Nguyen:18,Niculae:17} on the transcripts described in Section~\ref{sec:data}.
We noticed that since the two systems were trained to classify only claims and premises, they were never able to correctly predict warrants in our transcripts.
Additionally, both systems classified the overwhelming majority of moves as premise, resulting in negative kappa in some cases.
Using our  scheme to create a corpus of classroom discussion data manually annotated for argumentation, specificity, and knowledge domain will support the development of more robust NLP prediction systems.

\section{Conclusions}
In this work we proposed a new annotation scheme for three theoretically-motivated features of student talk in classroom discussion: argumentation, specificity, and knowledge domain.
We demonstrated usage of the scheme by presenting an annotated  excerpt of a classroom discussion. We demonstrated that the scheme can be annotated with high reliability and reported on scheme validity. Finally, we discussed some possible applications and challenges posed by the proposed annotation scheme for both the educational and NLP communities.
We plan to extend our annotation scheme to label information about collaborative relations between different argument moves, and release a corpus annotated with the extended scheme.

\section*{Acknowledgements}
We want to thank Haoran Zhang, Tazin Afrin, and Annika Swallen for their contribution, and all the anonymous reviewers for their helpful suggestions.

This work was supported by the Learning Research and Development Center at the University of Pittsburgh.


\bibliography{bea_2018_annotations}
\bibliographystyle{acl_natbib}

\end{document}